DIFFERENTIAL GEOMETRY/PDE SEMINAR

Monday, August 14 Padelford C-401 2:00 pm

Solutions of the Maxwell-Einstein Equations

Nina Zipser (MIT)

In this talk I will discuss the existence of unique, globally hyperbolic, and geodesically complete solutions of the Maxwell-Einstein equations which are “close” to Minkowski Space. This work is a generalization of Christodoulou and Klainerman’s proof of the nonlinear global stability of Minkowski Space. No previous understanding of General Relativity will be assumed. The talk will be geared towards an audience with some background in differential geometry or partial differential equations.